OpenStudy (anonymous):

no one can get the right answer..ive been wrong 3 times.. help please solve (1)symbolically (2) numerically 6x+3y= 6 -6x+2y=-6

OpenStudy (anonymous):

3y=6-6x y=2(1-x)

OpenStudy (anonymous):

for the first equation divide by 3 so you get 2x + y = 6, so then y = -2x + 6 and x = (-y+6)/2 this is symbolic

OpenStudy (anonymous):

-6x+2(2(1-x))=-6 -10x+4=-6 -10x=-10 x=1

OpenStudy (anonymous):

then sub x=1 into an equation